a container contains neon at 27 degrees Celsius. its temperature rises to 243 krad celsius .find the change in kinetic energy of neon atoms


1
Expert's answer
2021-01-31T18:57:48-0500

By the definition of the kinetic energy, we get:


"KE=\dfrac{3}{2}RT."

Then, we can write:


"\Delta KE=KE_f-KE_i=\dfrac{3}{2}R(T_f-T_i),""\Delta KE=\dfrac{3}{2}\cdot8.314\ \dfrac{J}{K\cdot mol}\cdot(516.15\ K-300.15\ K)=2694\ J."